# A Taxonomy of Questions for Critical Reflection in Machine-Assisted Decision-Making

### v2

@@ Abstract.
Decision-makers run the risk of relying too much on machine recommendations, which is associated with lower cognitive engagement. Reflection has been shown to increase cognitive engagement and improve critical thinking and reasoning and therefore decision-making. However, there is currently no approach to support reflection in machine-assisted decision-making. We therefore present a taxonomy that serves to systematically create questions related to machine-assisted decision-making that promote reflection and thus cognitive engagement and ultimately a deliberate decision-making process. Our taxonomy builds on a taxonomy of Socratic questions and a question bank for human-centred explainable AI (XAI), and illustrates how XAI techniques can be utilised and repurposed to formulate questions. As a use case, we focus on clinical decision-making. An evaluation in education confirms the applicability and expected benefits of our taxonomy. Our work contributes to the growing research on human-AI interaction that goes beyond the paradigm of machine recommendations and explanations and aims to enable effective human oversight as required by the European AI Act.
@@ 1. Introduction
Reflection, also referred to as reflective practice or reflective thought, is an important skill in professional practices (Schön, 1983). It is beneficial in, for example, management (Cunliffe, 2002), law (Casey, 2014), and healthcare (Mann et al., 2009; Sandars, 2009), and is part of many educational curricula (Boud and Walker, 1998). As the American philosopher John Dewey (1933, p.9) states, reflection consists of an
“active, persistent and careful consideration of any belief or supposed form of knowledge in light of the grounds that support it and further conclusion to which it tends”.
As reflection helps to scrutinise the validity of information and assumptions (Mezirow, 1990), it has the potential to improve reasoning, judgement, and problem-solving (Ghanizadeh, 2017; Khoshgoftar and Barkhordari-Sharifabad, 2023; Mamede and Schmidt, 2023). Consequently, reflection has been shown to improve strategic decisions (Walger et al., 2016) and diagnostic accuracy (Mamede et al., 2008; Prakash et al., 2019; Hess et al., 2015). In addition, reflection enables decision-makers to recognise and articulate the reasons for their decisions, which helps them to direct their actions accordingly and take responsibility (forward-looking) and accountability (backward-looking). Responsible decision-making ideally leads to more desirable and fairer outcomes for all parties involved. In view of these effects, it seems sensible to encourage and support reflection during decision-making, especially as many high-impact decisions nowadays are informed by machine recommendations.
Decision-support systems (DSS) assist decision-makers in domains like law, finance and healthcare. Physicians, for example, can use clinical DSSs to make a diagnosis or find a treatment option for a patient. Studies show, however, that operators tend to rely too much on these systems and accept incorrect recommendations (Dratsch et al., 2023), which is known as overreliance (Passi and Vorvoreanu, 2022). Moreover, explainable AI (XAI) (Longo et al., 2024), a common approach for calibrating reliance, has mixed and even negative effects, such as increasing overreliance (Bansal et al., 2021; Bussone et al., 2015; Ehsan et al., 2021). In light of this, and to reduce harmful consequences of overreliance, the European AI Act (European Parliament and Council of the European Union, 2024, §14) requires:
@@
In this paper, we propose a potential solution for mitigating overreliance, namely cognitive engagement of the decision-maker through questions that encourage critical reflection on the decision at hand, including the machine recommendation and the data that led to it. By critical reflection, we mean the above-mentioned practice of evaluating relevant information, without taking information at face value, and analysing assumptions that may have been taken for granted (Hong and Choi, 2011; Sengers et al., 2005). In order to elicit reflection, the decision-maker can be asked questions, for which the Socratic questioning method offers a helpful starting point. The Socratic method, named after the Greek philosopher Socrates, involves systematically asking questions to explore complex ideas, uncover assumptions, and gain a better understanding of the topic under investigation (Paul, 2019).
Although the Socratic method is slowly gaining traction in the field of human-machine interaction (Liu et al., 2024; Lara, 2021; Ang et al., 2023), there is as yet no systematic approach to applying this questioning method to the context of machine-assisted decision-making. To address this gap, we therefore provide a taxonomy of questions that serves the creation of questions for critical reflection in machine-assisted decision-making. For our taxonomy we synthesise prior work, namely 1) a taxonomy of Socratic questions (primarily for education) (Paul, 2019), 2) a question bank for the design of human-centred XAI (Liao et al., 2020, 2021; Liao and Varshney, 2022), and 3) a revised version of Bloom’s taxonomy for categorising educational objectives (Krathwohl, 2002; Forehand, 2010).
We primarily focus on clinical decision-making, where a physician uses a DSS to find a treatment option for patients. In this context, there may be multiple solutions with uncertain outcomes, so deliberate decision-making is crucial. An evaluation of the applicability of our taxonomy in the education domain shows however that our taxonomy can be applied to other domains without adaptation. Overall, our proposed taxonomy builds on evidence from literature that critical reflection can increase cognitive engagement of the decision-maker. This paper makes the following contributions:
@@
We present a new taxonomy to help formulate relevant questions that stimulate reflection in machine-assisted decision-making (section 4). This taxonomy supports the design and development of interventions that go beyond machine recommendations and explanations. In addition, our taxonomy can benefit other stakeholders such as developers, auditors, or consumers who need to remain critical in the development, evaluation, and use of DSS (section 6).
@@ 2. Background
In the following, we will discuss background work that led us to the creation of our question taxonomy. In order to sketch the problem of overreliance on machine recommendations and explanations (section 2.1), we begin with a short review of the relevant literature on human-machine interaction, particularly in relation to the human factor in explainable AI (section 2.2). From alternative interaction approaches, we derive that encouraging cognitive engagement is a promising method to reduce overreliance (section 2.3). We identify the benefits of reflective thought for cognitive engagement as it promotes critical thinking, reasoning, and thus responsible decision-making (section 2.4). Finally, we suggest that questions can be useful to stimulate reflection and thus cognitive engagement, and introduce the Socratic questioning method, which is widely used in education (section 2.5). In short, to arrive at our taxonomy, we combine relevant insights from empirical findings and literature on human-machine interaction, education, and philosophy.
2.1. Overreliance on DSS
@@
Decision-support systems (DSS) generally process information to arrive at an output, i.e., recommendation, by using (simple) mathematical models, or more advanced machine learning models, including deep learning approaches. A meta-study of 106 experimental studies on machine-assisted decision-making found that decision-making accuracy generally decreases when decision-makers use DSS (Vaccaro et al., 2024). Decision-makers may tend to prefer machine recommendations over their own judgement (Tschandl et al., 2020), even when the recommendations are wrong (Jacobs et al., 2021; van der Stigchel et al., 2023). The acceptance of incorrect recommendations is referred to as overreliance, which can be because of low vigilance or lack of cognitive engagement (Passi and Vorvoreanu, 2022).
@@
A common approach to mitigating overreliance is explainable AI (XAI) (Vasconcelos et al., 2023), which can, for example, provide information about which input was decisive for a particular output, i.e., recommendation. It is assumed that these explanations will help the operator to assess the machine recommendation and thus make an informed decision about whether or not to follow the recommendation. Provided explanations are, however, not always taken into account (van der Waa et al., 2021). Further, the aforementioned meta-study also found that explanations do not lead to significant synergy effects in terms of the performance of the human-machine team (Vaccaro et al., 2024). On the contrary, explanations may increase the likelihood of operators accepting a machine recommendation regardless of its correctness (Bansal et al., 2021; Bussone et al., 2015; Ehsan et al., 2021), thereby increasing overreliance (Vered et al., 2023). One reason for this is that the focus on numerical explanations can give a false sense of objectivity, leading operators, regardless of their expertise, to overestimate the capabilities of a DSS (Ehsan et al., 2024). Consequently, the provision of additional information in the form of explanations does not automatically lead to more cognitive engagement of the decision-maker and more accurate decisions (Gajos and Mamykina, 2022).
@@ 2.2. A Question Bank for XAI
To provide more actionable explanations based on user needs, a question bank for explainable AI helps designers and developers select the appropriate XAI technique to provide a relevant explanation that supports the operator’s goal, e.g., informed decision-making. To this end, the question bank maps prototypical user questions to available XAI techniques (Liao et al., 2020, 2021; Liao and Varshney, 2022). Feature importance, for example, can provide insights into why-questions, such as “Why was this prediction given?”.
Answering the (technical) question of how a DSS arrived at a particular output is however not always desired or helpful. As one participant working on clinical decision-support systems mentioned in the study on the XAI question bank:
“[explanations by system rationale] are essentially ‘this is how I do it, take it or leave it’. But doctors don’t like this approach…Thinking that [AI is] giving treatment recommendations is the wrong place to start, because doctors know how to do it. It’s everything that happens around that decision they need help with… more discussions about the output, rather than how you get there” (Liao et al., 2020, p.6).
While the XAI question bank allows for the design of more human-centred XAI, it is still very algorithmic- and data-centred, as we will show (Figure 1) and as the authors themselves note. Even if it is known how a DSS arrived at an output or recommendation, operators might be more interested in factors outside the model (Zednik, 2019; Ehsan and Riedl, 2024), which could relate to questions like, “What is missing from the dataset?”, “Are there any other overlooked alternative solutions?”, or “How to integrate this recommendation into the decision-process?”. Scholars therefore advocate moving from the current algorithmic focus of XAI methods to a broader, socio-technical perspective of explanations (Ehsan et al., 2021; Ehsan and Riedl, 2024; Ghassemi et al., 2021).
With our proposed taxonomy, we want to stimulate “more discussions about the output” (study participant in Liao et al., 2020) by identifying and categorising relevant question types. As we will show, current XAI techniques can be utilised and explanations can be repurposed to formulate and inform possible questions. Framing causal explanations as questions, for example, was found to improve human judgement (Danry et al., 2023). As such, in line with actionability, we consider “explanations as a means to help a data subject [and decision-maker] act rather than merely understand” (Wachter et al., 2017, p.843), where ‘act’ in this case is deliberate decision-making through reflection.
@@ 2.3. Increasing Cognitive Engagement
To mitigate overreliance, scholars explore different interactions with machine recommendations and explanations that attempt to cognitively engage the decision-maker in the decision-making process. We will briefly mention four relevant approaches: 1) promoting cognitive engagement through interventions, 2) supporting the decision-maker to make their own decision by presenting evidence for and against, 3) formulating explanations as questions, and 4) proposals to stimulate reflection.
@@
Second, instead of presenting a recommendation and justifying it with an explanation that the decision-maker must accept or reject, a so-called hypothesis-driven recommender shows evidence for and against a decision (Miller, 2023). In this way, the system supports the decision-maker’s reasoning process and gives them the control to make an informed decision themselves. A study found that while presenting evidence for and against did little to improve diagnostic accuracy, physicians valued the reflective aspect and had more confidence in their final decisions (Cabitza et al., 2023).
Third, in an experiment causal explanations were presented both as statements and questions (Danry et al., 2023). For example, the explanation as a statement: “The patient has the flu because of a headache”, is formulated as a question: “If the patient has a headache, does it follow that they have the flu?”. The study found that explanations framed as questions improved human judgement about the logical validity of the information provided. Similar to the hypothesis-driven recommender, questions help the decision-maker to think for themselves (Ryan and Deci, 2000), which leads to more cognitive engagement (Gajos and Mamykina, 2022).
Fourth, it has been suggested that DSS should take on roles other than that of a recommender, such as a devil’s advocate raising objections and challenging the decision-maker (Chiang et al., 2024; Sarkar, 2024; Ma et al., 2024b). Insights from decision-making in human groups show that a devil’s advocate has positive effects on decision performance (Schwenk, 1984; Schweiger et al., 1989). Similarly, others argue that DSS should support deliberation and facilitate reflection (Ma et al., 2024a; Zhang et al., 2023; Haselager et al., 2023; Schmidt and Mamede, 2023; Cornelissen et al., 2022).
In view of the above, we understand questions as cognitive interventions (1), which facilitate own reasoning and decision-making (2&3). With our taxonomy, we therefore aim to both build on and work towards these four approaches by offering more potential questions compared to 3) and concretising the proposals for promoting deliberation and reflection (4).
@@ 2.4. Reflection and Responsible Decision-Making
In areas where decisions have far-reaching consequences, such as law or healthcare, decision-makers have a professional responsibility; physicians, for example, are committed to the well-being of patients. This professional responsibility involves an epistemological responsibility, which means that decision-makers are responsible for gathering and evaluating relevant information, and knowing the reasons for a particular course of action (van Baalen and Boon, 2015). In regard to past actions, knowing and being able to provide reasons for an action allows for backward-looking responsibility or accountability. More importantly, however, knowing the reasons makes it possible to determine and direct future actions in order to achieve a desired state, and thus to take forward-looking responsibility (Loughran, 1996). Reflection-in-action can therefore increase professional responsibility, both in a backward-looking and, more importantly, in a forward-looking sense. Ideally, this forward-looking responsibility leads to fairer and better decision outcomes for all parties involved (Ueda et al., 2024). Reflection-in-action could, for example, prevent a physician from jumping to conclusions after seeing seemingly similar patient cases throughout the day, which could lead to tunnel vision (Mamede and Schmidt, 2023). Reflection makes it possible to distinguish between relevant and irrelevant information, and, to quote Dewey again, “converts action that is merely appetitive, blind and impulsive into intelligent action” (Dewey, 1933, p.17).
2.5. The Socratic Questioning Method

### v3

@@ Abstract.
Decision-makers run the risk of relying too much on machine recommendations, which is associated with lower cognitive engagement. Reflection has been shown to increase cognitive engagement and improve critical thinking and therefore decision-making. Questions are a means to stimulate reflection, but there is a research gap regarding the systematic creation and use of relevant questions for machine-assisted decision-making. We therefore present a taxonomy of questions aimed at promoting reflection and cognitive engagement in order to stimulate a deliberate decision-making process. Our taxonomy builds on the Socratic questioning method and a question bank for explainable AI. As a starting point, we focus on clinical decision-making. Brief discussions with two medical and three educational researchers provide feedback on the relevance and expected benefits of our taxonomy. Our work contributes to research on mitigating overreliance in human-AI interactions and aims to support effective human oversight as required by the European AI Act.
@@ 1. Introduction
Reflection, also referred to as reflective practice or reflective thought, is an important skill in professional practices (Schön, 1983). It is beneficial in, for example, management (Cunliffe, 2002), law (Casey, 2014), and healthcare (Mann et al., 2009), and is part of many educational curricula (Boud and Walker, 1998). As the American philosopher John Dewey (1933, p.9) states, reflection is an
“Active, persistent, and careful consideration of any belief or supposed form of knowledge in the light of the grounds that support it and further conclusion to which it tends”.
As reflection helps to scrutinise the validity of assumptions and information (Mezirow, 1990), it has the potential to improve reasoning, judgement, and problem-solving (Ghanizadeh, 2017; Khoshgoftar and Barkhordari-Sharifabad, 2023; Mamede and Schmidt, 2023). Consequently, reflection has been shown to improve strategic decisions (Walger et al., 2016) and diagnostic accuracy (Mamede et al., 2008; Hess et al., 2015; Prakash et al., 2019). In addition, reflection enables decision-makers to recognise and articulate the reasons for their decisions, allowing them to direct their actions accordingly and take responsibility (forward-looking) and accountability (backward-looking). Responsible decision-making ideally leads to more desirable outcomes for all parties involved. In view of this, it seems sensible to encourage and support reflection during decision-making, especially as many high-impact decisions nowadays are informed by machine recommendations or predictions.
Decision-support systems (DSS) provide recommendations that assist decision-makers in domains like law, finance and healthcare. Physicians, for example, can use clinical DSSs to make a diagnosis or find a treatment option. Studies show, however, that operators tend to rely too much on these systems and accept incorrect recommendations (Dratsch et al., 2023), which is known as overreliance (Passi and Vorvoreanu, 2022). In order to reduce harmful consequences of overreliance, the European AI Act (European Parliament and Council of the European Union, 2024, §14) requires:
@@
Various approaches therefore aim to mitigate overreliance by increasing the decision-maker’s cognitive engagement through different interactions, such as presenting evidence for and against a decision (Miller, 2023) or presenting explanations in the form of questions (Danry et al., 2023). Others argue that DSS should support deliberation and reflection rather than make recommendations (Schmidt and Mamede, 2023; Ma et al., 2024). Similarly, in previous work we have introduced the idea of a reflection machine, which asks questions to the decision-maker (Cornelissen et al., 2022; Haselager et al., 2023; Fischer, 2025). However, it remains unclear what kind of questions should be asked in order to effectively stimulate cognitive engagement and reflection.
To address this gap, we provide a taxonomy of questions that helps to systematically formulate relevant questions that stimulate critical reflection on the decision at hand, including the machine recommendation and the data that led to it. To identify relevant questions, the Socratic questioning method, named after the Greek philosopher Socrates, proves to be a useful starting point as it involves systematically asking questions to explore complex ideas, uncover assumptions, and gain a better understanding of the topic under investigation (Paul, 2019). As such, Socratic questioning has been shown to increase students’ critical thinking (Dalim, 2022; Ho et al., 2023). Although the Socratic method is slowly gaining traction in the field of human-machine interaction (Lara, 2021; Ang et al., 2023; Liu et al., 2024), there is as yet no systematic approach for applying this questioning method to the context of machine-assisted decision-making.
We therefore adapt a taxonomy of Socratic questions based on concepts of critical thinking (Paul, 2019) to identify aspects to which questions should relate, and transfer it to the context of machine-assisted decision-making by synthesising it with a question bank for the design of human-centred explainable AI (Liao et al., 2020, 2021; Liao and Varshney, 2022). We also draw inspiration, albeit only marginally, from a revised version of Bloom’s taxonomy (Krathwohl, 2002; Forehand, 2010) to identify cognitive processes to which questions can refer.
While we aim for a domain-independent taxonomy by building on general concepts of critical thinking, we take as a starting point clinical decision-making, where a physician uses a DSS to find a treatment option among several possibilities with uncertain outcomes for patients. To assess the transferability of our taxonomy, we apply it to the educational domain, where teachers use learning analytics dashboards to make decisions about students. We have discussed our taxonomy with medical and educational researchers, who have emphasised the relevance and expected benefits of our taxonomy. Nevertheless, an empirical evaluation of the effectiveness of certain questions in a specific decision-making context remains for future work. The contributions of this paper are as follows:
@@
We present a new taxonomy to help formulate relevant questions that stimulate reflection in machine-assisted decision-making (section 4). This taxonomy supports the design and development of interventions that go beyond machine recommendations and explanations. In addition, our taxonomy can benefit stakeholders such as developers, auditors, or consumers who need to remain critical in the development, evaluation, and use of DSS (section 6).
@@ 2. Background
In the following, we will discuss background work that led us to the creation of our question taxonomy. In order to sketch the problem of overreliance on machine recommendations and explanations (section 2.1), we begin with a short review of the relevant literature on human-machine interaction, particularly in relation to the human factor in explainable AI (section 2.2). From alternative interaction approaches, we derive that promoting cognitive engagement is a promising method to reduce overreliance (section 2.3). We identify the benefits of reflective thought for cognitive engagement as it promotes critical thinking, reasoning, and thus responsible decision-making (section 2.4). Finally, we suggest that questions can be useful to stimulate reflection and thus cognitive engagement, and introduce the Socratic questioning method, which is widely used in education (section 2.5). In short, to arrive at our taxonomy, we combine relevant insights from empirical findings and literature on human-machine interaction, education, and philosophy.
2.1. Overreliance on DSS and Explainable AI
@@
Decision-support systems (DSS) provide recommendations based on information by using (simple) mathematical models, or more advanced machine learning models, including deep learning approaches. A meta-study of 106 experimental studies on machine-assisted decision-making found that decision-making accuracy generally decreases when decision-makers use DSS (Vaccaro et al., 2024). Decision-makers may tend to prefer machine recommendations over their own judgement (Tschandl et al., 2020), even when the recommendations are wrong (Jacobs et al., 2021). The acceptance of incorrect recommendations, referred to as overreliance, can be because of low vigilance or lack of cognitive engagement (Passi and Vorvoreanu, 2022).
A common approach to mitigating overreliance is explainable AI (XAI) (Vasconcelos et al., 2023; Longo et al., 2024), which can, for example, provide information about which input was decisive for a particular output. It is assumed that these explanations will help the operator to assess the DSS recommendation and thus make an informed decision as to whether or not to follow the recommendation.
@@
The aforementioned meta-study, however, found that explanations do not lead to significant synergy effects in human-machine decision-making (Vaccaro et al., 2024). On the contrary, explanations may increase the likelihood of operators accepting a machine recommendation regardless of its correctness (Bansal et al., 2021; Bussone et al., 2015; Ehsan et al., 2021), thereby increasing overreliance (Vered et al., 2023). One reason for this is that the focus on numerical explanations can give a false sense of objectivity, leading operators, regardless of their expertise, to overestimate the capabilities of a DSS (Ehsan et al., 2024). Consequently, providing additional information in the form of explanations does not automatically lead to more cognitive engagement of the decision-maker and to more accurate decisions (Gajos and Mamykina, 2022).
@@ 2.2. A Question Bank for XAI
To provide more relevant and actionable explanations, a question bank for explainable AI helps designers and developers select the appropriate XAI technique that supports the operator’s information needs (Liao et al., 2020, 2021; Liao and Varshney, 2022). To this end, the question bank clusters and maps prototypical user questions to XAI techniques. Feature importance, for example, can provide insights into why-questions, such as “Why was this prediction given?”.
Answering the (technical) question of how a DSS arrived at a particular output is however not always desired or helpful. As one participant working on clinical decision-support systems mentioned in the study on the XAI question bank (Liao et al., 2020, p.6):
[explanations by system rationale] are essentially ‘this is how I do it, take it or leave it’. But doctors don’t like this approach…Thinking that [AI is] giving treatment recommendations is the wrong place to start, because doctors know how to do it. It’s everything that happens around that decision they need help with… more discussions about the output, rather than how you get there.
Although the XAI question bank allows for the design of more human-centred XAI, it is still very algorithmic- and data-centred, as the authors themselves state and as we will discuss below. Even if it is known how a DSS arrived at an output or recommendation, operators might be more interested in factors regarding the broader socio-technical context outside of the model (Zednik, 2019; Ehsan et al., 2021; Ghassemi et al., 2021; Ehsan and Riedl, 2024), which could relate to questions like, “What is missing from the dataset?”, “Are there any other overlooked alternative solutions?”, or “How to integrate this recommendation into the decision-process?”.
With our proposed taxonomy, we want to stimulate “more discussions about the output” (study participant in Liao et al., 2020) by identifying and categorising relevant questions. The XAI question bank will prove useful as it contains questions of operators and maps them to XAI techniques. As we will show, we utilise the mapping from questions to XAI techniques and reverse the process to repurpose explanations to formulate possible questions, e.g., “Is it relevant to focus on factor ?”. As such, in line with actionability, we consider “explanations [informing questions] as a means to help a data subject [and decision-maker] act rather than merely understand” (Wachter et al., 2017, p.843), where ‘act’ in this case is deliberate decision-making through reflection.
@@ 2.3. Increasing Cognitive Engagement
To mitigate overreliance, scholars explore different interactions with machine recommendations and explanations in addition to XAI that attempt to cognitively engage the decision-maker in the decision-making process. We will briefly mention four relevant approaches: 1) promoting cognitive engagement through interventions, 2) supporting the decision-maker to make their own decision, 3) formulating explanations as questions, and 4) proposals to stimulate reflection.
@@
Second, instead of presenting a recommendation and justifying it with an explanation that the decision-maker must accept or reject, a so-called hypothesis-driven recommender shows evidence for and against a decision (Miller, 2023). In this way, the system supports the decision-maker’s reasoning process and allows them to make an informed decision themselves. A study found that while presenting evidence for and against did little to improve diagnostic accuracy, physicians valued the reflective aspect and had more confidence in their final decisions (Cabitza et al., 2023).
Third, one study found that explanations formulated as questions improve human judgement of the logical validity of the information provided (Danry et al., 2023). Causal explanations were presented both as statements and questions. For example, the statement: “The patient has the flu because of a headache”, is formulated as a question: “If the patient has a headache, does it follow that they have the flu?”. Similar to the hypothesis-driven recommender, questions help the decision-maker to think for themselves (Ryan and Deci, 2000), which leads to more cognitive engagement (Gajos and Mamykina, 2022).
Fourth, it has been suggested that DSS should take on roles other than that of a recommender, such as a devil’s advocate raising objections and challenging the decision-maker (Chiang et al., 2024; Ma et al., 2024; Sarkar, 2024). Insights from decision-making in human groups show that a devil’s advocate has positive effects on decision performance (Schwenk, 1984; Schweiger et al., 1989). Similarly, others argue that DSS should support deliberation and facilitate reflection (Cornelissen et al., 2022; Haselager et al., 2023; Schmidt and Mamede, 2023; Zhang et al., 2023; Ma et al., 2025).
In view of the above, we understand questions as cognitive interventions (1), which facilitate own reasoning and decision-making (2&3). With our taxonomy, we therefore want to both build on and expand these four approaches by offering more potential questions compared to (3) and supporting the design of suggested interactions that promote reflection (4).
@@ 2.4. Reflection and Responsible Decision-Making
In areas where decisions have far-reaching consequences, such as law or healthcare, decision-makers have great professional responsibility. Physicians, for example, are committed to the well-being of patients. This professional responsibility includes an epistemological responsibility that requires decision-makers to gather and evaluate relevant information, and to know the reasons for a particular course of action (van Baalen and Boon, 2015). Being able to provide reasons for (past) actions allows for backward-looking responsibility or accountability. More importantly, knowing the reasons enables one to determine and direct future actions to achieve a desired state, and thus to take forward-looking responsibility (Loughran, 1996).
Reflection-in-action can therefore increase epistemological and thus professional responsibility, both in a backward-looking and, more importantly, in a forward-looking sense. Ideally, this forward-looking responsibility leads to fairer and more desirable decision outcomes for all parties involved (Ueda et al., 2024). Reflection-in-action could, for example, prevent a physician from jumping to conclusions after seeing seemingly similar patient cases throughout the day, which could lead to tunnel vision (Mamede and Schmidt, 2023). Reflection makes it possible to distinguish between relevant and irrelevant information, and, to quote Dewey (1933, p.17) again, “converts action that is merely appetitive, blind and impulsive into intelligent action”.
2.5. Types of Questions